Exercise 6.14 Let y be distributed Bernoulli P(y = 1) unknown 0<p<1 p and P(y = 0) = 1-p f or Some

(a) Show that p E( (b) Write down the natural moment estimator p of . (c) Find var (p) (d) Find the asymptotic distribution of vn (-p) as no. as n> OO.
Answer #1

Part 1

mathbb{E}(Y)=sum y imesmathbb{P}(Y=y)=0*(1-p)+1*p=p

Part 2

Suppose we have iid observations 1, 12; ...1 n from the distribution of Y.

Define ar{Y}_n=rac1nsum Y_i .

Then, mathbb{E}(ar{Y}_n)=rac1n summathbb{E}(Y_i)=rac1nsum p=p .

So, the natural moment estimator hat{p}=ar{Y}_n

Part 3

Let us first derive the variance of Y.

Var(Y)=mathbb{E}(Y^2)-(mathbb{E}(Y))^2=(0^2*(1-p)+1^2*p)-(p)^2=p-p^2=p(1-p)

Now, p(1-p) Var(þ) = Var(%) =-Var(y) C7

Part 4

Let T=sqrt{n}(hat{p}-p) .

Then, T=sqrt{n}(ar{Y}_n-p)=sqrt{n}(rac1nsum Y_i-p)=sqrt{n} imesrac1nsum(Y_i-p) .

Now, (Y_i-p)s are iid random variables, with mathbb{E}(Y_i-p)=mathbb{E}(Y_i)-p=p-p=0 and Var(Y_i-p)=Var(Y_i)=p(1-p) .

Hence, by Lindeberg Levy Central Limit Theorem, asymptotic distribution of T is N0,(1-p)) .